Q. A rotating ball hits a rough horizontal plane with a vertical velocity v and angular velocity w.Given that the coefficient of friction is u and the vertical component of the velocity after the collision is v/2.find


a)the angular velocity after collision[ans.w-(15vu/4R)


b.)the impulsive ground reaction during the collision.{ans :3mv/2 * root of 1+u^2}

Grade:11

2 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Raj, 

Let J1 be the horizontal impulse and J2 be the vertical impulse. J1=uJ2, J2=mv/2, let angular velocity after collision=w2, angular impulse=RJ1=uRJ2=umvR/2=(2mR2/5)(w-w2),5uv=4R(w-w2), w2=w-(5uv/4R), J=√(J12+J22)=(mv/2)√(1+u2)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras

jayant sharma
13 years ago

as told by chetan nayakar , the method is correct but the vertical impulse is 3mv/2 not mv/2  .  so multiply his answer by 3 and you will get your answer